# High-dimensional Laplace asymptotics up to the concentration threshold

### v1

@@ 6.1 Preliminaries on cumulants
When there is repetition among the $Y_{j}$, there is an alternative expression for the cumulant. Let $\alpha=(\alpha_{1},\dots,\alpha_{M})$ with $\alpha_{j}\geq 0$. Suppose that $Y_{1},\dots,Y_{m}$ consist of $\alpha_{1}$ copies of $X_{1}$, $\alpha_{2}$ copies of $X_{2}$, and so on, up to $\alpha_{M}$ copies of $X_{M}$. Let $|\alpha|=\alpha_{1}+\dots+\alpha_{M}=m$. Then
@@
To prove Lemma 4.3, we need to show
@@
6.2 The term $b_{1}$
Let us use (6.8) with $M=2$ to compute $b_{1}(f,g)$. Only $\alpha=(2,0)$ and $\alpha=(0,1)$ satisfy $\alpha_{1}+2\alpha_{2}=2$. Thus
@@ 6.2 The term $b_{1}$
Then $p_{1}(x)=T_{1}[x]+T_{3}[x^{\otimes 3}]$ and $p_{2}(x)=T_{2}[x^{\otimes 2}]+T_{4}[x^{\otimes 4}]$. The polynomial $p_{1}$ is odd, therefore $\mathbb{E}[p_{1}(Z)]=0$ and $\mathrm{Var}(p_{1}(Z))=\mathbb{E}[p_{1}(Z)^{2}]$. We now rewrite $p_{1}$ using Hermite polynomials. For $x=(x_{1},\dots,x_{d})$ define the first- and third-order multivariate Hermite polynomials [5, Chapter 1.3]
@@
The Hermite polynomials are orthogonal in $L^{2}(\mathcal{N}(0,I_{d}))$ [5, Chapter 1.3]. In other words, if $Z\sim\mathcal{N}(0,I_{d})$, then $\mathbb{E}[H_{i}(Z)H_{j}(Z)]=0$ if $i\neq j$, $\mathbb{E}[H_{i}(Z)H_{jk\ell}(Z)]=0$ for any $i,j,k,\ell$, and $\mathbb{E}[H_{ijk}(Z)H_{\ell mn}(Z)]=0$ if $(i,j,k)\neq(\ell,m,n)$, viewed as unordered triplets. Furthermore, we have $\mathbb{E}[H(Z_{i})^{2}]=1$ and it is straightforward to show $\mathbb{E}[H_{ijk}(Z)^{2}]=1=6/3!$ if $i,j,k$ are distinct, $\mathbb{E}[H_{iij}(Z)^{2}]=2=6/3$ if $i,j$ are distinct, and $\mathbb{E}[H_{iii}(Z)^{2}]=6=6/1$. Thus for general $i,j,k$, the expectation $\mathbb{E}[H_{ijk}(Z)^{2}]$ is given by $6$ divided by the number of distinct ways to rearrange the indices $i,j,k$.
@@
Since cumulants are multilinear, and since the coefficients of the $p_{k}$’s are entries of the derivative tensors $F_{k+2}$, which are uniformly bounded, it suffices to show that $|\mathrm{cum}(X(t)^{b_{1}},\dots,X(t)^{b_{n}})-\mathrm{cum}(Z^{b_{1}},\dots,Z^{b_{n}})|=o(t^{M})$ for $n$ arbitrary multi-indices $b_{i}=(b_{i}^{1},\dots,b_{i}^{d})$, $i=1,\dots,n$. Here, $Z^{b_{i}}=Z_{1}^{b_{i}^{1}}\dots Z_{d}^{b_{i}^{d}}$, and similarly for $X(t)^{b_{i}}$. But $\mathrm{cum}(X(t)^{b_{1}},\dots,X(t)^{b_{n}})$ is a polynomial function of moments $\mathbb{E}[X(t)^{b}]$ [31, (3.2.11)]. Thus it suffices to bound $|\mathbb{E}[X(t)^{b}]-\mathbb{E}[Z^{b}]|$ for any multi-index $b$. We have

### v2

@@ 6.1 Preliminaries on cumulants
When there is repetition among the $Y_{j}$, there is an alternative expression for the cumulant. Specifically, suppose we have another set of random variables $X_{k}$, $k=1,2,3,\dots$. Let $\alpha=(\alpha_{1},\dots,\alpha_{M})$ with $\alpha_{j}\geq 0$. Suppose that $Y_{1},\dots,Y_{m}$ consist of $\alpha_{1}$ copies of $X_{1}$, $\alpha_{2}$ copies of $X_{2}$, and so on, up to $\alpha_{M}$ copies of $X_{M}$. Let $|\alpha|=\alpha_{1}+\dots+\alpha_{M}=m$. Then
@@
For convenience, we recall the formula in Lemma 4.3, which we will prove in Section 6.3.
@@
6.2 More on the terms $b_{k}$
We first study $b_{M/2}$ to determine the highest-order $f$ derivative contributing to it. This is useful in Section 8 below, where we study the derivative order of $b_{L-1}(f,g)-b_{L-1}(f,1)$. We then compute $b_{1}(f,g)$ explicitly.
By (6.4), the highest-order $f$ derivatives in the formula (6.8) for $b_{M/2}$ necessarily arise from $\alpha$ such that $\alpha_{M}>0$. But since $\sum_{i=1}^{M}i\alpha_{i}=M$, if $\alpha_{M}>0$ then we must have $\alpha_{M}=1$ and $\alpha_{i}=0$ for all $i\neq M$. Thus the highest-order $f$ derivative contribution to $b_{M/2}(f,g)$ is
$\begin{split}\frac{\mathrm{cum}(p_{0,\dots,0,1}(Z))}{M!}&=\frac{\mathrm{cum}(p_{M}(Z))}{M!}=\frac{\mathbb{E}[p_{M}(Z)]}{M!}\\ &=\frac{\mathbb{E}[\nabla^{M}(\log g)(0)[Z^{\otimes M}]]}{M!}-\frac{\mathbb{E}[\nabla^{M+2}f(0)[Z^{\otimes M+2}]]}{(M+2)!}.\end{split}$ (6.9)
Next, we use (6.8) with $M=2$ to compute $b_{1}(f,g)$. Only $\alpha=(2,0)$ and $\alpha=(0,1)$ satisfy $\alpha_{1}+2\alpha_{2}=2$. Thus
@@ 6.2 More on the terms $b_{k}$
Then $p_{1}(x)=T_{1}[x]+T_{3}[x^{\otimes 3}]$ and $p_{2}(x)=T_{2}[x^{\otimes 2}]+T_{4}[x^{\otimes 4}]$. The polynomial $p_{1}$ is odd, therefore $\mathbb{E}[p_{1}(Z)]=0$ and $\mathrm{Var}(p_{1}(Z))=\mathbb{E}[p_{1}(Z)^{2}]$. We now rewrite $p_{1}$ using Hermite polynomials. For $x=(x_{1},\dots,x_{d})$ define the first- and third-order multivariate Hermite polynomials [6, Chapter 1.3]
@@
The Hermite polynomials are orthogonal in $L^{2}(\mathcal{N}(0,I_{d}))$ [6, Chapter 1.3]. In other words, if $Z\sim\mathcal{N}(0,I_{d})$, then $\mathbb{E}[H_{i}(Z)H_{j}(Z)]=0$ if $i\neq j$, $\mathbb{E}[H_{i}(Z)H_{jk\ell}(Z)]=0$ for any $i,j,k,\ell$, and $\mathbb{E}[H_{ijk}(Z)H_{\ell mn}(Z)]=0$ if $(i,j,k)\neq(\ell,m,n)$, viewed as unordered triplets. Furthermore, we have $\mathbb{E}[H(Z_{i})^{2}]=1$ and it is straightforward to show $\mathbb{E}[H_{ijk}(Z)^{2}]=1=6/3!$ if $i,j,k$ are distinct, $\mathbb{E}[H_{iij}(Z)^{2}]=2=6/3$ if $i,j$ are distinct, and $\mathbb{E}[H_{iii}(Z)^{2}]=6=6/1$. Thus for general $i,j,k$, the expectation $\mathbb{E}[H_{ijk}(Z)^{2}]$ is given by $6$ divided by the number of distinct ways to rearrange the indices $i,j,k$.
@@
Since cumulants are multilinear, and since the coefficients of the $p_{k}$’s are entries of the derivative tensors $F_{k+2}$, which are uniformly bounded, it suffices to show that $|\mathrm{cum}(X(t)^{b_{1}},\dots,X(t)^{b_{n}})-\mathrm{cum}(Z^{b_{1}},\dots,Z^{b_{n}})|=o(t^{M})$ for $n$ arbitrary multi-indices $b_{i}=(b_{i}^{1},\dots,b_{i}^{d})$, $i=1,\dots,n$. Here, $Z^{b_{i}}=Z_{1}^{b_{i}^{1}}\dots Z_{d}^{b_{i}^{d}}$, and similarly for $X(t)^{b_{i}}$. But $\mathrm{cum}(X(t)^{b_{1}},\dots,X(t)^{b_{n}})$ is a polynomial function of moments $\mathbb{E}[X(t)^{b}]$ [39, (3.2.11)]. Thus it suffices to bound $|\mathbb{E}[X(t)^{b}]-\mathbb{E}[Z^{b}]|$ for any multi-index $b$. We have